OpenStudy (anonymous):

, a canoeist travels 48 miles at a certain speed. The canoeist travels 14 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 4 hrs. What was the speed on each part of the trip?

OpenStudy (anonymous):

48/x+14/(x-5)=4 (48(x-5)+14x)/x(x-5)=4 48x-240+14x=4(x^2-5x) 62x-240=4x^2-20x 0=4x^2-82x+240 use quad. form. and that will give you x.

OpenStudy (anonymous):

awesome, thank you so very much... I have been working on this problem for about 35 minutes.. UGH! thank you very much for your help!!!

OpenStudy (anonymous):

do you understand how i got the initial equation?

OpenStudy (anonymous):

yes, I am pretty sure that I do... by entering the time and miles already traveled...

OpenStudy (anonymous):

total distance divided by speed (dist/time) will give you time

OpenStudy (anonymous):

ok, the word problems are getting complicated and so frustrating...

OpenStudy (anonymous):

its a matter of looking at one part at a time and figuring out what it is really saying.

OpenStudy (anonymous):

yes, I think I got it, now... I will come back if I need more help. Thank you for taking the time...

OpenStudy (anonymous):

you are welcome. this is my way of killing time between classes.

OpenStudy (anonymous):

I got it... thank you.. 1st part of the trip was 16.96 and the second part was 11.96 YOU ROCK!!! THANK YOU